# Accessing proton number fluctuations in limited regions of phase space

### v1

@@ IV Back to back covariance for protons
The emission of particles in opposite directions can be correlated due to the collective flow and conservation of the total transverses momentum. On the other hand, local fluctuations in the baryon density in the fireball are not expected to significantly correlate the emission back to back (Fig. 2). The covariance of the proton number in two back to back azimuthal angle intervals is expected to be dominated by the contribution from volume fluctuations. Such volume fluctuations can be estimated using measurements for reference particles - charged mesons.
Figure 6: The covariance of the number of protons in two back to back azimuthal angle windows as a function of the window width. The results without additional fluctuations in the baryon density are represented with the solid line and the results with additional fluctuations are represented with the dotted line. The dashed and dash-dotted lines represent the estimates of the contribution of volume fluctuations to the covariance (Eq. 25).
@@
To estimate the volume fluctuation contribution for the covariance of protons emitted back to back let us consider the joint probability distribution to observe a number of protons $N_{0}$, $N_{\pi}$ and mesons $M_{0}$, $M_{\pi}$ in two azimuthal angle intervals centered at $\phi=0$ and $\pi$
@@
is a 4-dimensional Gaussian distribution for the emission volumes of protons and mesons in the two azimuthal angle intervals. Here,
@@
$\Sigma_{4}=\begin{pmatrix}\sigma_{\cal N}^{2}&Cov_{{\cal N},{\cal M}}&Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N},{\cal M}}&\sigma_{\cal M}^{2}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&\sigma_{\cal M}^{2}&Cov_{{\cal N},{\cal M}}\\ Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}&Cov_{{\cal N},{\cal M}}&\sigma_{\cal N}^{2}\end{pmatrix}\ .$ (21)
The upper left and lower right $2\times 2$ submatrices match the covariance matrix for the two-dimensional Gaussian distribution in Eq. (9). There are three additional covariances between emission volumes in opposite azimuthal angle intervals
$Cov_{{\cal N}_{0},{\cal N}_{\pi}}$ $=$ $\langle N_{0}N_{\pi}\rangle-\langle N\rangle^{2}=Cov_{bb}^{proton}$
@@
$Cov_{bb}^{vol}=\frac{-2Cov(p,m)Cov_{bb}(p,m)C_{2}^{meson}-Cov_{bb}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}$ (25)
@@
$C_{2}^{vol,4D}=\frac{-2Cov_{bb}^{meson}Cov_{bb}(p,m)Cov_{bb}(p,m)+C_{2}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}\ .$ (26)
I have verified that the expression (26) for the volume fluctuations contribution to the proton factorial cumulant yields results numerically very close to the simpler expression in Eq. (16). On the other hand, the expression (25) will be used in the following as an estimate of the volume fluctuation contribution to the covariance of proton multiplicities observed in two back to back azimuthal angle windows.
@@
The covariance of proton multiplicities observed at opposite azimuthal angles is nearly independent on the strength of local fluctuations in the deposition of baryons, for $\Delta\phi\leq\pi$ (compare solid and dash-dotted lines in Fig. 6). The proton covariance closely matches the contribution from volume fluctuations (dashed and dotted lines in Fig. 6). In narrow azimuthal angle windows ($\Delta\phi=\pi/4$) the covariance of proton multiplicities is almost entirely explained by volume fluctuations; even for the most extreme baryon deposition fluctuations assumed (Fig. 7).
As the strength of baryon density fluctuations increases, the second factorial cumulant of the proton number increases beyond the volume fluctuation contribution (Fig. 4), while the covariance of proton multiplicities in two azimuthal angle intervals remains approximately constant (Fig. 7). Consequently, the dynamical correlation coefficient (defined with the second factorial cumulant in the denominator)
@@
V Summary and Outlook
The paper presents a method to search for local fluctuations in the fireball produced in relativistic heavy-ion collisions. In particular, I study the fluctuations of a specific type of particles of interest (protons) with respect to reference particles (charged mesons). The method allows to partly separate global fluctuations of the particle emission from fluctuations on smaller scales. In order to be sensitive to local fluctuations in the transverse plane, I propose to consider fluctuations of particle emission in restricted windows of azimuthal angle. Due to the presence of a strong transverse flow in heavy-ion collisions, particles emitted at a given azimuthal angle originate predominantly from a region of the fireball located at that same angle. Strong global fluctuations in the particle emission occur due to fluctuations of the collective flow event by event, e.g. fluctuations of the magnitude and orientation of the harmonic flow vectors. Such global volume fluctuation contributions can be estimated calculating fluctuations of the meson distribution and scaling accordingly.
@@ V Summary and Outlook
In particular, the method discussed here could be used in the search for fluctuations in baryon density in the vicinity of the QCD critical point [28, 19]. In order to validate the method based on the study of proton distributions in restricted intervals of azimuthal angle further studies are planned. These involve : the calculation for net baryons and net protons, using a model with nontrivial higher order proton cumulants, including effects of global baryon conservation, including effects of baryon diffusion, and making predictions for different collisions energies using realistic, 3+1-dimensional hydrodynamic simulations. Finally, the dependence of the baryon factorial cumulants on the width of the azimuthal angle window could be potentially sensitive to the value of the correlations length of the underlying baryon fluctuations.
@@ Acknowledgments
The author thanks Adam Bzdak for useful discussions. This research was supported by the AGH University of Science and Technology and by the Polish National Science Centre grant: 2023/51/B/ST2/01625.

### v2

@@ IV Back to back covariance for protons
The emission of particles in opposite directions can be correlated due to the collective flow and conservation of the total transverse momentum. On the other hand, local fluctua- tions of the baryon density in the fireball are not expected to induce significant correlations in the multiplicities of protons emitted back to back (Fig. 2). The covariance of the proton number in two back to back azimuthal angle intervals is expected to be dominated by the contribution from volume fluctuations. Such volume fluctuations can be estimated using measurements for reference particles - charged mesons.
Figure 6: The covariance of the number of protons in two back to back azimuthal angle windows as a function of the window width. The results without additional fluctuations in the baryon density are represented with the solid line and the results with additional fluctuations are represented with the dotted line. The dashed and dash-dotted lines represent the respective estimates of the contribution of volume fluctuations to the covariance (Eq. 27).
@@
To estimate the volume fluctuation contribution for the covariance of protons emitted back to back one can consider the joint probability distribution to observe a number of protons $N_{0}$, $N_{\pi}$ and mesons $M_{0}$, $M_{\pi}$ in two azimuthal angle intervals centered at $\phi=0$ and $\pi$
@@
is a four-dimensional Gaussian distribution for the emission volumes of protons and mesons in the two azimuthal angle intervals. Here,
@@
$\Sigma_{4}=\begin{pmatrix}\sigma_{\cal N}^{2}&Cov_{{\cal N},{\cal M}}&Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N},{\cal M}}&\sigma_{\cal M}^{2}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&\sigma_{\cal M}^{2}&Cov_{{\cal N},{\cal M}}\\ Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}&Cov_{{\cal N},{\cal M}}&\sigma_{\cal N}^{2}\end{pmatrix}\$. (23)
The upper left and lower right $2\times 2$ submatrices match the covariance matrix for the two-dimensional Gaussian distribution in Eq. (9). There are three additional covariances between emission volumes in opposite azimuthal angle intervals:
$Cov_{{\cal N}_{0},{\cal N}_{\pi}}$ $=$ $\langle N_{0}N_{\pi}\rangle-\langle N\rangle^{2}=Cov_{bb}$
@@
$Cov_{bb}^{vol}=\frac{2Cov(p,m)Cov_{bb}(p,m)C_{2}^{meson}-Cov_{bb}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}$ (27)
@@
$C_{2}^{vol,4D}=\frac{-2Cov_{bb}^{meson}Cov_{bb}(p,m)Cov(p,m)+C_{2}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}\ .$ (28)
I have verified that the expression (28) for the volume fluctuation contribution to the proton factorial cumulant yields results numerically very close to the simpler expression in Eq. (16). On the other hand, the expression (27) will be used in the following as an estimate of the volume fluctuation contribution to the covariance of proton multiplicities observed in two back to back azimuthal angle windows.
@@
The covariance of proton multiplicities observed at opposite azimuthal angles is nearly independent of the strength of local fluctuations in the deposition of baryons, for $\Delta\phi\leq\pi$ (compare solid and dash-dotted lines in Fig. 6). The proton covariance closely matches the contribution from volume fluctuations (dashed and dotted lines in Fig. 6). In narrow azimuthal angle windows ($\Delta\phi=\pi/4$) the covariance of proton multiplicities is almost entirely explained by volume fluctuations; even for the most extreme baryon deposition fluctuations assumed (Fig. 7).
As the strength of baryon density fluctuations increases, the second factorial cumulant of the proton number increases beyond the volume fluctuation contribution (Fig. 4), while the covariance of proton multiplicities in two opposite azimuthal angle intervals remains approximately constant (Fig. 7). Consequently, the dynamical correlation coefficient (defined with the second factorial cumulant in the denominator)
@@
V Conclusions
The paper presents a method to search for local fluctuations in the fireball produced in relativistic heavy-ion collisions. In particular, I study the fluctuations of a specific type of particles of interest (protons) with respect to reference particles (charged mesons). The method allows one to partly separate global fluctuations of the particle emission from fluctuations on smaller scales. In order to be sensitive to local fluctuations in the transverse plane, I propose to consider fluctuations of particle emission in restricted windows of azimuthal angle. Due to the presence of a strong transverse flow in heavy-ion collisions, particles emitted at a given azimuthal angle originate predominantly from a region of the fireball located at that same angle. Strong global fluctuations in particle emission occur due to fluctuations of the collective flow event by event, e.g., fluctuations of the magnitude and orientation of the harmonic flow vectors. Such global volume fluctuation contributions can be estimated calculating fluctuations of meson distributions and scaling accordingly.
@@ V Conclusions
In particular, the method discussed here could be used in the search for fluctuations of the baryon density in the vicinity of the QCD critical point [28, 19]. In order to validate the method based on the study of proton distributions in restricted intervals of azimuthal angle further studies are planned. These involve the calculation for net baryons and net protons, using a model with nontrivial higher order proton cumulants, including effects of global baryon conservation, including effects of baryon diffusion, and making predictions for different collisions energies using realistic, 3+1-dimensional hydrodynamic simulations. Using a full 3+1-dimensional dynamics with baryon diffusion could moderate the final effect of baryon density fluctuations. Finally, the dependence of the baryon factorial cumulants on the width of the azimuthal angle window could be potentially sensitive to the value of the correlation length of the underlying baryon fluctuations.
@@ Acknowledgments
The author thanks Adam Bzdak for useful discussions. This research was supported by the AGH University of Science and Technology and by the Polish National Science Centre Grant No. 2023/51/B/ST2/01625.
Data availability
This manuscript has associated data openly available at [51].